# Training-free Camera Control for Video Generation

### v3

@@ 4 Experiments
Figure 3: Qualitative comparisons with finetuned methods. CamTrol’s outputs align well with complex trajectories from reference video, while others fail to perceive subtle camera pose changes. We provide videos in the supplementary materials for a clearer comparison.
@@
SVD MotionCtrl CameraCtrl CamTrol ($t_{0}=10$)
@@
Figure 4: Generalization comparisons. CamTrol can avoid domain collapse that rises from overfitting on certain datasets, adapting to more general scenes (Left, where input does not resemble real scenes), and maintain video’s dynamics while adhering to desired camera movements (Right).
@@ 4.1 Experimental Settings
We compare our method with state-of-the-art works including MotionCtrl (Wang et al., 2023) and CameraCtrl (He et al., 2024). To ensure a fair comparison, we employ SVD (Blattmann et al., 2023a) as base model for all methods. SVD is originally trained on resolution of $576\times 1024$, but SVD-based CameraCtrl only support $320\times 576$. Since changing the original resolution leads to suboptimal generation quality, we use $576\times 1024$ for MotionCtrl and CamTrol, then resize their outputs into $320\times 576$ for calculating metrics. For all methods, the number of frames and the decoding size of SVD are set to 14. We use 25 steps for both inversion and generation processes.
@@
In quantitative evaluation, FVD (Unterthiner et al., 2018), FID (Heusel et al., 2017) and IS (Saito et al., 2020) are used to assess video generation quality, and CLIPSIM (Wu et al., 2021) quantifies the similarity between generated video and input prompt. For the accuracy of camera motion, we adopt ParticleSFM(Zhao et al., 2022) and produce estimated camera trajectories from generated videos, with the use of Absolute Trajectory Error(ATE) measuring their differences compared to ground truth. Relative Pose Error (RPE) is calculated to assess between consecutive frames how well the relative motions match expected ones including their transition (RPE-T) and rotation part (RPE-R). Specifically, we randomly sample 500 prompt-trajectory pairs from RealEstate10k (Zhou et al., 2018), and use the corresponding videos as reference for calculating FVD and FID. Since SVD is an image-to-video model, we generate the first frames using Stable Diffusion (Rombach et al., 2022) by text prompts. We also compare the results produces by vanilla SVD as a reference. For camera motion accuracy, we provide the evaluations on ground truth videos as a lower bound of these metrics.
@@ 4.2 Comparisons with State-of-the-art Methods
Figure 5: Comparison with base model. Controlling camera motion via prompt engineering doesn’t work at most times. Instead, CamTrol offers robust control towards video’s camera movement in a training-free manner.
Figure 6: Effectiveness of layout prior. Layout prior guidance compensates for the vacancies (Left) and static contents (Right) caused by point cloud rendering.
@@
Quantitative evaluations are shown in Table  1. In the table, the best results are in bold, and the second best is underlined. The performance of vanilla SVD (without motion control) is indicated as SVD, while the lower bound of motion metrics provided by ground truth videos is indicated as Reference. In terms of video quality, CamTrol attains comparable performance with methods finetuned on the RealEstate10k dataset. For motion accuracy, CamTrol also gets lowset score in ATE and RPE-T/R. The quantitative evaluation demonstrates CamTrol’s ability to generate videos with both accurate camera motion and high visual quality.
@@
Qualitative comparisons are illustrated in Fig. 3. The reference trajectory includes zoom, pan and roll. While MotionCtrl and CameraCtrl fail to perceive subtle camera motions, resulting in simple pan movement, CamTrol is able to follow the complex trajectory and generate video with correct move. We also evaluate the generalization ability of different methods in generating more general scenes and dynamic content, the results are shown in Fig. 4. Since both MotionCtrl and CameraCtrl are finetuned on limited scenes (i.e., real estate videos) with static content, they can hardly generalize to other scenes, such as videos in non-realistic style or videos that include people. As illustrate in Fig. 4, their motion controls in both situations are misaligned with input movements. Furthermore, finetuning on such datasets leads to the loss of dynamics, which is a crucial element in video generation. In comparison, CamTrol preserves most of the prior knowledge from video base models, enabling it to handle general scenes as well as generate dynamic content. Relevant videos are available in the supplementary materials.
@@
We provide the computational analysis in Table 2, including the maximum GPU memory required and the time consumption for all methods during inference process. The evaluations are conducted under unified settings. We test on $576\times 320$, with 25 generation steps, the number of frames and decoding size of SVD are set to 14. As a training-free method, CamTrol requires no extra GPU memory during inference process compared with base model. This makes it save 10-20GB of GPU memory than other methods under the same circumstances, and makes it applicable running on a single RTX 3090. The major consumption of CamTrol comes from the time rendering multi-view images. Since this process is currently sequential in time, i.e., $t_{0}\to t_{1}\to t_{2}$, a more parallel approach may be more time-efficient. We will leave this to future works. The results of $576\times 1024$ resolution and more detailed settings can be found in the Appendix.
@@ 4.3 Ablation Study
$t_{0}=5$ 1079.88 68.52 7.14 0.3100 4.17 1.09 0.012
$t_{0}=10$ 778.46 68.06 7.05 0.3110 4.17 1.07 0.010
$t_{0}=15$ 754.14 67.98 7.00 0.3107 4.13 1.02 0.008
Figure 7: Effect of $t_{0}$. Smaller $t_{0}$ encourages
dynamics while larger $t_{0}$ preserves camera
@@
Figure 8: 3D rotation videos at different motion scales. CamTrol supports camera movements over various scales.
@@
To demonstrate that changes of camera motion are attributed to our method rather than the innate capability of video model, we conduct ablation study to assess its effectiveness. We add prompts describing certain camera moves (e.g. zooms out), letting video model understand by itself. The results are shown in Fig. 5. It could be observed that even provided with prompts indicating how camera should move, base model fail to produce correct results. Instead, CamTrol  is able to implement designated motion control without any instructions from text prompts. In Table 1, the comparison with vanilla SVD also demonstrate CamTrol’s effectiveness.
@@
We employ ablation study to validate the effectiveness of layout prior guidance, illustrating its necessity from two aspects: the completeness of vacancies and dynamic of generated video. In Fig. 6, we showcase frames before and after noise prior guidance. With camera pose changes, there appears regions unfilled in point cloud and causes blank spaces in rendered images(left part); Besides, due to the static nature of point cloud, rendered images remain stationary(right part). Noise layout prior could compensate for these flaws, finally produce videos with inpainted vacancies and rationalized dynamics.
@@
$t_{0}$ is a crucial factor that influences the trade-off between generated video’s diversity and its faithfulness to camera motions requirements. To investigate its effect on the output, we conduct experiments with various $t_{0}$ values, relevant results are shown in Fig. 8. As illustrated, videos generated with larger $t_{0}$ tend to conform better to camera motion requirements, but suffer from decrease in dynamics; On the contrary, smaller $t_{0}$ leads to more plausible generations but fails to meet camera’s requirements, as latents at these timesteps carry more randomness. We also provide quantitative evaluations of different $t_{0}$ in Table 3. As $t_{0}$ increase, CamTrol produces videos alike to static, camera-moving scenes (which have lower FVD as we use RealEstate10k as reference videos) with higher accuracy in motion control.
@@
Our proposed CamTrol can be seamlessly plugged and played under various scenarios, accommodating different base models, resolutions, generating length, all in training-free style. We present visual results of its applications under different settings, including CogVideoX (Yang et al., 2024b)(diffusion transformer model, $720\times 480$ resolution, 49 frames) and VideoFusion (Luo et al., 2023)(decomposed diffusion process, $128\times 128$ resolution, 16 frames), in Fig. 10. Our approach remains effective applied to alternative video base models, resolutions and generating lengths, demonstrating its strong robustness and generalization ability.
@@ 4.4 Further Applications
Figure 9: Multi-trajectory video generation. CamTrol is able to generate videos with different trajectories for one scene.
Figure 10: Applied onto CogVideoX (Yang et al., 2024b) and VideoFusion (Luo et al., 2023). CamTrol can be plugged and played under various situations, accommodating different base models, different resolutions, different generating lengths, all in a training-free manner.
@@
One of the most advantages of our method is it can generate videos with rotating moves and produce outputs similar to 3D generation models (Voleti et al., 2024; Melas-Kyriazi et al., 2024). While these 3D models need large-scale training on 3D dataset and could only handle inputs in specific styles, our approach is able to deal with any type of images and achieve this in a completely zero-shot manner. One example of this is shown in Fig. 8.
@@
By combining different basic camera trajectories, CamTrol can support hybrid camera movements and endow generated video with cinematic charm. Besides this, explicit motion modeling also equips CamTrol with the ability to support trajectories containing precise extrinsics, and generate videos presenting any complicated camera movement. Videos avaiable at https://lifedecoder.github.io/CamTrol/.
@@
One natural application of our method is to generate multi-trajectory videos for one scene. Since the point cloud is fixed once the reconstruction process finished, the consistency of the content between different trajectories is guaranteed, and we can easily produce multiple camera-moving videos of this scene. We present this application in Fig. 9. More results on multi-trajectory video generation could be found in the supplementary materials.
@@
CamTrol supports camera movements at controllable scales. By specifying different magnitudes of camera’s extrinsic matrix within point cloud spaces, rendered images will exhibit varying degrees of motion, leading to videos with distinct scales of camera movements. This further demonstrate the powerful controllablity of CamTrol, and provides a new pathway for video’s customized camera control. We provide relevant visualization in Fig. 8.
@@ 5 Conclusion
In this paper, we propose a training-free and robust method CamTrol to offer camera control for off-the-shelf video diffusion models. It consists of two-stage procedure including explicit camera motion modeling in 3D point cloud space and video generation utilizing layout prior of noisy latents. Compared to previous work, CamTrol  does not require any additional finetuning on camera-annotated datasets or self-supervised training via data augmentation, simultaneously, comprehensive experiments demonstrate its superior performance in both generation video quality and motion alignment against other state-of-the-art methods. Furthermore, we show the ability of CamTrol  generalizing to various scenarios, as well as its impressive applications including unsupervised generation of 3D rotation video, scalable motion control and dealing with complicated camera trajectories.
@@ Appendix A More Results on Camera Control
We showcase additional qualitative results of CamTrol on various camera movements. As pdf format is not the best way to exhibit videos, we put them on our demo page:
If not accessible, please refer to the supplementary materials for a quicker view to the demo page.
This demo page includes CamTrol-generated videos including basic camera motions (including Zoom, Tilt, Pan, Pedestal, Truck, Roll, Rotate, Hybrid, Complicated, detailed definitions are in B), hybrid motions (Zoom In first, then Pedestal Up, Zoom Out + Pedestal Up + Truck Left + Tilt Down + Pan Right) and complicated motions generated from precise camera extrinsics (extracted from RealEstate10k (Zhou et al., 2018)).
Besides, it contains 3D rotation videos generated unsupervisedly from video base models (both objects and scenes). These outputs share sort of similarity with outputs of 3D generation models, as they all exhibit in a turning-table like way, which camera rotates around some objects. The difference here is that 3D model, as trained on specific datasets, could only generate outputs in certain styles, e.g., single static object with no background. Instead, our model could handle arbitrary image as input, and generate a rotating-around video with proper dynamics. From this aspect, our method could be seen as a infinite source of attaining 3D data. And by utilizing our method with stronger backbones, video foundation models could truly become the largest source of 3D data as it should be.
Furthermore, it contains additional results mentioned in ablation studies, e.g. controlling camera motions at different scales, and the effectiveness of using layout prior compared with the raw output given by video base model.
@@ Appendix B Definitions of Basic Camera Motions
We refer to the terminology in cinematography to describe different camera motions, definitions of each type are detailed in Table 5.
To get consistent images from multiple views, we set camera motion as a trajectory of extrinsic matrix $\left\{\mathbf{P}_{1},...,\mathbf{P}_{N-1}\right\}$, each including a rotation matrix and translation matrix representing camera’s pose and position. For hybrid motions, CamTrol supports both spatial (i.e. moving several basic moves simultaneously) and temporal (concatenation of basic motions sequentially) combinations. In the case of complicated trajectories, one can directly use precise parameters for camera’s extrinsic matrix as input to control video’s motion. Additionally, if these parameters not available, a reference video with corresponding move can also serve as input. With the help of SFM algorithms (e.g. COLMAP  [^1] https://github.com/colmap/colmap, ParticleSFM (Zhao et al., 2022)), camera motion can be easily estimated and used for imitation. In this sense, CamTrol can also be seen as an unsupervised method to achieve video motion transfer.

### v4

@@ 4 Experiments
Figure 3: Qualitative comparisons with finetuned methods. CamTrol’s outputs align well with complex trajectories from the reference video, while others fail to capture subtle changes in camera pose. We provide videos in the supplementary materials for clearer comparison.
@@
SVD MotionCtrl CameraCtrl CamTrol ($t_{0}=15$)
@@
Figure 4: Generalization comparisons. CamTrol can avoid domain collapse that arises from overfitting on certain datasets and adapt to more general scenes (Left), meanwhile preserving video’s dynamics while adhering to desired camera movements (Right).
@@ 4.1 Experimental Settings
We compare our method with state-of-the-art works including MotionCtrl (Wang et al., 2023) and CameraCtrl (He et al., 2024). To ensure a fair comparison, we employ SVD (Blattmann et al., 2023a) as base model for all methods. SVD is originally trained at a resolution of $576\times 1024$, but SVD-based CameraCtrl only supports $320\times 576$. Since changing the original resolution results in suboptimal generation quality, we use $576\times 1024$ for MotionCtrl and CamTrol, then resize the outputs to $320\times 576$ to calculate metrics. For all methods, the number of frames and the decoding size of SVD are set to 14. We use 25 steps for both the inversion and generation processes.
@@
In the quantitative evaluation, FVD (Unterthiner et al., 2018), FID (Heusel et al., 2017) and IS (Saito et al., 2020) are used to assess video generation quality, and CLIPSIM (Wu et al., 2021) quantifies the similarity between the generated video and the input prompt. For camera motion accuracy, we adopt ParticleSFM(Zhao et al., 2022) to estimate camera trajectories from generated videos, with the use of Absolute Trajectory Error (ATE) to measure their differences from the ground truth. Relative Pose Error (RPE) is calculated to assess how well the relative motions between consecutive frames match the expected ones, including their translation (RPE-T) and rotation component (RPE-R). Specifically, we randomly sample 500 prompt-trajectory pairs from RealEstate10k (Zhou et al., 2018), and use them as references for calculating FVD and FID. Since SVD is an image-to-video model, we generate the first frames using Stable Diffusion (Rombach et al., 2022) based on text prompts. We also provide the results produced by vanilla SVD as a reference. For camera motion accuracy, we provide the evaluations on ground truth videos as a lower bound of these metrics.
@@ 4.2 Comparisons with State-of-the-art Methods
Figure 5: Comparison with base model. Controlling camera motion via prompt engineering rarely succeeds. Instead, CamTrol offers robust control to video’s camera movements.
Figure 6: Effectiveness of layout prior. Layout prior guidance compensates for the voids (Left) and static elements (Right) in point cloud.
@@
Quantitative evaluations are shown in Table  1. In the table, the best results are in bold, and the second best are underlined. The performance of vanilla SVD (without motion control) is denoted as SVD, while the lower bound for motion metrics, provided by ground truth videos, is denoted as Reference. In terms of video quality, CamTrol attains performance comparable to methods finetuned on the RealEstate10k dataset. For motion accuracy, CamTrol also achieves the lowest score in ATE and RPE-T/R. The quantitative evaluation demonstrates CamTrol’s ability to generate videos with both accurate camera motion and high visual quality.
@@
Qualitative comparisons are illustrated in Fig. 3. The reference trajectory includes zoom, pan, and roll. While MotionCtrl and CameraCtrl fail to capture subtle camera motions, resulting in simple pan movements, CamTrol is able to follow the complex trajectory and generate videos with correct motion. We also evaluate the generalizability of different methods in generating more general scenes and dynamic content. The results are shown in Fig. 4. Since both MotionCtrl and CameraCtrl are finetuned on limited scenes (i.e., real estate videos) with static content, they have difficulty generalizing to other scenes, such as videos in non-realistic styles or videos that include people. As illustrated in Fig. 4, their motion controls in both situations are misaligned with the intended movements. Furthermore, finetuning on such datasets leads to a loss of dynamics, which is a crucial element in video generation. In comparison, CamTrol preserves most of the prior knowledge from video base models, enabling it to handle general scenes as well as generate dynamic content. Relevant videos are available in the supplementary materials.
@@
We provide the computational analysis in Table 2, including the maximum GPU memory required and the processing time for all methods during inference. Evaluations are conducted under unified settings. We test at a resolution of $576\times 320$ with 25 generation steps. The number of frames and the decoding size of SVD are set to 14. As a training-free method, CamTrol requires no extra GPU memory during the inference process compared to the base model. This saves 10-20GB of GPU memory compared to other methods under the same circumstances, allowing it to run on a single RTX 3090. The major consumption of CamTrol comes from rendering multi-view images. Since this process is temporally sequential, i.e., $t_{0}\to t_{1}\to t_{2}$, a more parallelized approach may improve time efficiency. The results for $576\times 1024$ resolution and more detailed settings can be found in the Appendix.
@@ 4.3 Ablation Study
$t_{0}=20$ 1079.88 68.52 7.14 0.3100 4.17 1.09 0.012
$t_{0}=15$ 778.46 68.06 7.05 0.3110 4.17 1.07 0.010
$t_{0}=10$ 754.14 67.98 7.00 0.3107 4.13 1.02 0.008
Figure 7: Effect of $t_{0}$. Larger $t_{0}$ encourages
dynamics while smaller $t_{0}$ preserves camera
@@
Figure 8: 3D rotation videos at different motion scales. CamTrol supports camera movements across multiple scales.
@@
To demonstrate that the changes in camera motion are attributed to our method rather than the innate capability of the video model, we conduct an ablation study to evaluate its effectiveness. We add prompts describing specific camera movements (e.g. zooms out) and let the video model interpret them on its own. The results are presented in Fig. 5. It can be observed that, even when provided with prompts indicating how the camera should move, the base model fails to produce correct results. In contrast, CamTrol is able to implement the designated motion control without any instructions from text prompts. In Table 1, the comparisons with vanilla SVD also demonstrate CamTrol’s effectiveness.
@@
We conduct an ablation study to validate the effectiveness of layout prior guidance, demonstrating its necessity in two aspects: the completeness of missing areas and the dynamics of the generated video. In Fig. 6, we showcase frames before and after applying noise prior guidance. As camera pose changes, there are regions left unfilled in the point cloud, leading to blank spaces in rendered images (left part); Additionally, due to the static nature of point cloud, the rendered images remain stationary (right part). Layout prior guidance compensates for these flaws, ultimately producing videos with inpainted gaps and natural dynamics.
@@
$t_{0}$ is a crucial factor that influences the trade-off between generated video’s diversity and its faithfulness to camera motion requirements. To investigate its effect on the output, we conduct experiments with various $t_{0}$ values, the relevant results are shown in Fig. 8. As illustrated, videos generated with smaller $t_{0}$ tend to conform better to camera motion requirements, but suffer from decreased dynamics; On the other hand, larger $t_{0}$ leads to more plausible generations but fails to meet camera’s requirements, since the latents at these timesteps carry more randomness. We also provide quantitative evaluations for different $t_{0}$ values in Table 3. As $t_{0}$ decreases, CamTrol produces videos resembling static, camera-moving scenes (which have lower FVD as we use RealEstate10k as the reference videos) with higher accuracy in motion control.
@@
Our proposed CamTrol can be seamlessly integrated into various scenarios, accommodating different base models, resolutions, and generation lengths, all in a training-free style. We present visual results of its applications under different settings, including CogVideoX (Yang et al., 2024b) (diffusion transformer model, $720\times 480$ resolution, 49 frames) and VideoFusion (Luo et al., 2023) (decomposed diffusion process, $128\times 128$ resolution, 16 frames), in Fig. 10. Our approach remains effective when applied to alternative video base models, resolutions, and generation lengths, demonstrating its strong robustness and generalizability.
@@ 4.4 Further Applications
Figure 9: Multi-trajectory video generation. CamTrol is able to generate videos with different camera motions for the same scene.
Figure 10: Applied to CogVideoX (Yang et al., 2024b) and VideoFusion (Luo et al., 2023). CamTrol can be plug-and-play under various situations, accommodating different base models, different resolutions, different generation lengths, all in a training-free manner.
@@
One of the main advantages of our method is that it can generate videos with rotating movements and produce outputs similar to those of 3D generation models (Voleti et al., 2024; Melas-Kyriazi et al., 2024). While these 3D models require large-scale training on 3D datasets and can only handle inputs in specific styles, our approach is capable of dealing with any type of image and achieving this in a completely zero-shot manner. An example is shown in Fig. 8.
@@
By combining basic camera trajectories, CamTrol can support hybrid camera movements and produce videos with cinematic appeal. In addition, explicit motion modeling equips CamTrol with the ability to handle trajectories of precise extrinsics, allowing it to generate videos with arbitrarily complicated camera movements. Relevant results can be found at https://lifedecoder.github.io/CamTrol/.
@@
One natural application of our method is to generate multi-trajectory videos for a single scene. Since the point cloud remains fixed once the reconstruction has finished, the spatial consistency between different trajectories is guaranteed, making it easy to produce multiple camera-moving videos of the same scene. We demonstrate this application in Fig. 9. More results on multi-trajectory video generation can be found in the supplementary materials.
@@
CamTrol supports camera movements at controllable scales. By specifying different magnitudes of the camera’s extrinsic matrix, the rendered images will exhibit varying degrees of motion, resulting in videos with distinct scales of camera movements. This further demonstrates the powerful controllability of CamTrol and provides a new pathway for customized camera control of video generation. We provide relevant visualization in Fig. 8.
@@ 5 Conclusion
In this paper, we propose a training-free and robust method CamTrol to offer camera control for off-the-shelf video diffusion models. It consists of a two-stage procedure, including explicit camera motion modeling in 3D point cloud space and video generation utilizing the layout prior of noisy latents. Compared to previous works, CamTrol does not require additional finetuning on camera-annotated datasets or self-supervised training via data augmentation. Comprehensive experiments demonstrate its superior performance in both generation quality and motion alignment against other state-of-the-art methods. Furthermore, we show the ability of CamTrol to generalize to various scenarios, as well as its impressive applications including unsupervised generation of 3D video, scalable motion control, and dealing with complicated camera trajectories.
@@ Appendix A More Results on Camera Control
We present additional qualitative results of CamTrol on our demo page:
This demo page includes CamTrol-generated videos including basic camera motions (including Zoom, Tilt, Pan, Pedestal, Truck, Roll, Rotate, Hybrid, Complicated, detailed definitions are in Appendix B), hybrid motions (Zoom In first, then Pedestal Up, Zoom Out + Pedestal Up + Truck Left + Tilt Down + Pan Right), and complicated motions generated from precise camera extrinsics (extracted from RealEstate10k (Zhou et al., 2018)).
In addition, it contains 3D rotation videos generated unsupervised from video base models (both objects and scenes). These outputs share similarities with those of 3D generation models, as they all exhibit in a turning-table-like way, with the camera rotating around objects. The difference here is that the 3D model, as trained on specific datasets, can only generate outputs in certain styles, e.g., single static object with no background. Instead, our model can handle arbitrary images as input and generate rotating videos with appropriate dynamics. From this perspective, our method could be seen as an infinite source of 3D data. By employing our method with stronger backbones, video foundation models could truly become the largest source of 3D data, as they should be.
Furthermore, it contains additional results discussed in ablation studies, e.g., controlling camera motions at different scales, and the effectiveness of using layout prior compared with the raw output given by the video base model.
@@ Appendix B Definitions of Basic Camera Motions
We refer to the terminology in cinematography to describe different camera motions, the definitions of each type are detailed in Table 5.
To obtain consistent images from multiple views, we set camera motion as a trajectory of extrinsic matrices $\left\{\mathbf{P}_{1},...,\mathbf{P}_{N-1}\right\}$, each consisting of a rotation matrix and a translation matrix, representing the camera pose and position. For hybrid motions, CamTrol supports both spatial (i.e., performing several basic motions simultaneously) and temporal (i.e., concatenating basic motions sequentially) combinations. In the case of complicated trajectories, one can directly use precise parameters for camera’s extrinsic matrix as input to control video’s motion. Additionally, if these parameters are not available, a reference video with corresponding move can also serve as input. With the help of SFM algorithms (e.g., COLMAP  [^1] https://github.com/colmap/colmap, ParticleSFM (Zhao et al., 2022)), camera motion can be easily estimated and used for imitation. In this sense, CamTrol can also be seen as an unsupervised method to achieve video motion transfer.